OpenStudy (anonymous):

b^2 times b^5 without exponents is? Help please? I'm trying to find the formula on Google but I can't seem to find it. Not so that it shows me how to simplify without exponents. Thank you. :)

OpenStudy (mosaic):

b^2 times b^5 = b^(2+5) = b^7. Not sure how something with an exponent can be written without an exponent unless you want to use logarithm.

OpenStudy (anonymous):

I suppose that's what they want. It has to be without an exponent. I know that you just add exponents when they have the same base but this is asking me for the answer without exponents.

OpenStudy (mosaic):

Question does not make sense. Can you post a screenshot of the question?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Voila

OpenStudy (mosaic):

But in the lower part under "Fill in the blank" they say \(b^7\) which uses the exponent.

OpenStudy (anonymous):

Write I wrote that answer it asked me for it with and without exponents

OpenStudy (anonymous):

Right*

OpenStudy (mosaic):

Just doesn't make sense. Do you have any example in the book or online lessons that explains what they mean by writing without an exponent? Do they mean to write it as b7 instead of \(b^7\)?

OpenStudy (anonymous):

I don't think that's what they mean and no I don't have access to an online lesson...this is part of a practice quiz and while I'm taking this it locks me out of the lessons on my class...but I don't remember learning this anyway

OpenStudy (anonymous):

can you check my answer for another question in the meantime?

OpenStudy (mosaic):

ok. I can do one and then I have to log out.

OpenStudy (anonymous):

Thank you

OpenStudy (mosaic):

Try just "b" without the exponent "7" and see if it accepts the answer with the exponent omitted.

OpenStudy (anonymous):